Weight of oxygen in Fe2​O3​ and FeO is in the ratio of :

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
Dear student

One mole of Fe₂O₃ means 160 gram of Fe₂O₃

In 160 gram of Fe₂O₃ , weight of Oxygen = 16 × 3 = 48

Also, one mole of FeO means 72 gram of FeO

In 72 gram of FeO, weight of Oxygen = 16

Thus, the ratio of weight of Oxygen in Fe₂O₃  and FeO

= 48/16

= 3/1

3 : 1

thanks

team askiitians